# Human-in-the-Loop Optimization with Model-Informed Priors

### v1

@@ Abstract.
$c$s: Human-centered computing Interaction techniquesccs: Computing methodologies Machine learning
Figure 1. The general steps of our HOMI framework. Step 1: Model selection Designers or developers select a model, parameterized by $\theta$, that are relevant to the target task, and optionally fit the model’s parameters (e.g., Fitts’ law, with parameters $a$ and $b$) using minimal user data to better reflect the target context. Step 2: Synthetic user generation — A diverse set of synthetic users is created by sampling different parameter settings $\theta_{1},\theta_{2},\theta_{3},\dots$} from the fitted model’s parameter distribution. Step 3: Meta-BO training — The meta-optimizer interacts extensively with these synthetic users to learn efficient strategies for optimizing across user variability. Step 4: Deployment — The trained meta-BO is deployed with real users and quickly adapts to individual performance or preferences using the learned prior experience.
@@ 1. Introduction
Efficient and effective interface optimization for different users and contexts is a long-standing goal in human-computer interaction (HCI) (Norcio and Stanley, 1989; Zhai et al., 2002; Rao, 2019; Lavie and Meyer, 2010). Model-based optimization approaches use computational optimizers to search for interface designs that maximize predicted user performance, based on established models. The challenge of this approach is that practical models are scarce, and they are not able to capture specific users or contexts without model parameter inference (Beaumont et al., 2002; Moon et al., 2023), a known challenging step; therefore, this approach can not scale up easily or achieve true personalization. In recent years, human-in-the-loop optimization (HILO) emerged for its higher adaptability and flexibility (Lee et al., 2020; Khajah et al., 2016; Chiu et al., 2020; Kim et al., 2017). Unlike model-based methods, HILO operates directly on actual users: the optimizer proposes a design candidate, users interact with it, their performance is evaluated, and the optimizer updates the design candidate based on a series of observations. Various computational techniques, such as Bayesian optimization (Koyama et al., 2017; Liao et al., 2023; Brochu et al., 2010; Dudley et al., 2019), evolutionary and genetic algorithms (Battiti and Passerini, 2010; Meng and Wu, 2008), and heuristic-based techniques (Lee et al., 2020) have been explored in human-in-the-loop optimization. Among these, Bayesian optimization (BO) stands out due to its better performance and minimal assumptions about the task (Frazier, 2018). HILO with BO has been successfully applied in domains such as novel interaction techniques (Liao et al., 2020), design tools (Khajah et al., 2016; Brochu et al., 2010), and wearable interfaces (Liao et al., 2024a; Kim et al., 2017). Despite its adaptability, HILO often suffers from slow convergence due to a lack of prior information; it typically begins with a random search, requiring many iterations to find a good solution (Wang et al., 2022; Pu et al., 2024). This “cold-start” problem limits its deployment in real-time applications or dynamic environments.
@@
Recent work has attempted to mitigate this limitation by incorporating prior data from past users to initialize or guide the optimization process, a concept known as transfer learning or meta-learning. Prior data was used for example to construct a prior surrogate model (Liao et al., 2024b), to construct a series of prior models and aggregate them via a weighted-sum approach (Liao et al., 2024a), or, most recently to construct an adaptive surrogate model using Bayesian neural networks (Liao et al., 2025) on which the HILO approach was then trained before deployment to end-users. These works demonstrated that augmenting HILO with prior knowledge is a viable path toward solving the cold-start issue. However, a key bottleneck remains: the reliance on real human data. Prior transfer- or meta-learning–based methods require data from at least 10 users to be effective (Liao et al., 2024b, a), which can be prohibitively expensive or impractical. Moreover, any changes to the interaction task or interface require fresh rounds of data collection, severely limiting scalability. This motivates us to ask: Is it possible to train computational optimizers for higher online efficiency, yet, without relying on real human data?
@@
In this work, we propose a new HILO framework called Human-in-the-Loop Optimization with Model-Informed Priors (HOMI). Rather than relying on real user data for training the optimizer, HOMI leverages data from synthetic users simulated from predictive user models. This training phase allows Bayesian optimization to develop a broad understanding of the design space and learn effective adaptation strategies for diverse user contexts before the online-optimization process starts. A key advantage of using model-informed priors is that the optimizer can be trained with a wide variety of synthetic users, enabling generalization across different usage scenarios. In contrast, training with real users limits the optimizer’s knowledge to a small number of specific individuals, reducing its adaptability. At deployment, the trained optimizer switches to a human-in-the-loop mode, integrating prior knowledge with live user feedback to efficiently optimize the interface design in real time.
@@
We further demonstrate the effectiveness of our HOMI by applying NAF+ to real-time adaptation of a mid-air keyboard for direct touch input, which is an essential VR text entry technique, in a study with real users. The training involved synthetic users constructed via established models in HCI, including Fitts’ law (MacKenzie, 1992) and typing error model (Bi and Zhai, 2016). The results showed HOMI enabled by NAF+ outperforms the established baselines, including both manual keyboard setting and standard BO in that it significantly reduced the number of iterations required to personalize the keyboard.
@@ 2. Background & Related work
We review the key topics that are relevant to this work, especially focusing on (1) traditional optimization in HCI, (2) human-in-the-loop optimization (HILO), and the (3) techniques that augment Bayesian optimization with prior data for efficient interface adaptation.
@@ 2.1. Model-based Optimization in HCI
Optimization is an essential process in HCI and design, even if they are not explicitly articulated as such. In traditional user-centered design workflows, optimization is performed manually: designers generate a set of candidate designs, which are then evaluated through user testing and iteration (Abras et al., 2004). The field of computational interaction (Oulasvirta et al., 2018) aimed to integrate computational optimization methods in the design process to automatically identify optimal design candidates. This idea is not new: engineering fields have long used optimization workflows in which a solver proposes candidate designs, and these are evaluated through domain-specific models (e.g., physical, mechanical, chemical) to identify optimal solutions (Martins and Ning, 2021). This approach has proven successful in fields such as mechanical engineering (Rao et al., 2011), chemical engineering (Bhaskar et al., 2000), and architectural design (Michalek et al., 2002). HCI researchers have adopted similar methods: deploying computational solvers over user models to identify optimal interaction designs. Problems are often framed as assignment problems and solved using integer programming. Examples include menu optimization (Bailly et al., 2013; Todi et al., 2021), keyboard layout design (Feit et al., 2021; Karrenbauer and Oulasvirta, 2014), and UI layout adaptation in AR/VR environments (Lindlbauer et al., 2019; Evangelista Belo et al., 2022), and more. While this model-based approach has shown success in generating promising designs, it suffers from two key limitations: First, the application scope is highly limited by the availability of user models. Yet, practical models are scarce in real-world applications. Secondly, and more importantly, the user models represent an abstract “average user” but not any specific user under a specific context. Once deployed, such systems do not update or adapt their underlying model based on live observations, limiting their ability to adapt designs to individual needs or preferences. This work builds upon the foundation of model-based optimization but extends it into the domain of human-in-the-loop adaptation. In the HOMI framework, user models are not treated as fixed sources for optimization, but rather as generators of synthetic user data to train an optimizer. This repositions models from being optimization targets to being training resources, enabling scalable pre-training optimizers for HILO.
@@ 2.2. Human-in-the-Loop Bayesian Optimization
Human-in-the-loop optimization (HILO) is an emerging optimization paradigm in HCI that does not rely on predefined user models. Instead, HILO learns and optimizes directly from online observations, which may include user performance, preferences, or other forms of interaction feedback. The objective function $f$ is usually a user-centered performance metric (e.g., completion time, preferences, etc) that we want to optimize for (Liao et al., 2020, 2024a; Kim et al., 2017; Yamamoto et al., 2022). Each unique design candidate $x$ within the designated design space $\mathcal{X}$ results in a user-related objective function value, $f(x)$. The goal of HiLO is to identify the global optimal design $x^{*}$ that maximizes (or minimizes) the objective value $f(x^{*})$. Such optimization problem is challenging since the only way to access the user-relevant objective function is through costly user evaluations. Bayesian optimization (BO) is a popular computational method for human-in-the-loop optimization due to its minimum requirements and assumptions of the problem and reliable performances (Mockus, 1975; Shahriari et al., 2016). Within Bayesian optimization, a surrogate model, typically a Gaussian Process regression (GP), captures the properties of the target function. This model is continually refined with online function observations (the result of user evaluation), enhancing its prediction accuracy. An acquisition function utilizes the information of the surrogate model to computes the acquisition value (i.e., the worth or potential value) of a new candidate design. The candidate with the highest acquisition value is then selected for user evaluation in the subsequent iteration.
@@
Bayesian optimization has found many human-involved applications where the sample efficiency is particularly important, such as exoskeleton systems (Kim et al., 2017, 2019; Tucker et al., 2020), wearable interfaces (Liao et al., 2023; Guo et al., 2019), haptic displays (Liao et al., 2023; Catkin and Patoglu, 2023), input techniques (Liao et al., 2024b), and design tools (Brochu et al., 2010; Koyama et al., 2017; Mo et al., 2024). Despite its robust performance and relatively better efficiency over other optimization approaches (simulated annealing, evolutionary algorithms, etc.), a key limitation of BO lies in its “cold-start” problem. Since BO generally starts without any prior knowledge of the objective function, it must rely on random exploration in the early stages, which can slow down convergence. To address this, recent work has explored incorporating transfer learning and meta-learning to accelerate the optimization process, which we will review in the next subsection.
2.3. Transfer- and Meta-learning for BO
Transfer learning (Zhuang et al., 2020) and meta-learning (Wang, 2021) are similar concepts, both referring to enhancing the model training efficiency by leveraging the data gathered from previous similar tasks. Despite their differences in deep learning, these two terms are generally interchangeable in the context of Bayesian optimization. Bai et al. (2023) provides a review of recent advancements in this topic. In the following content we will refer to the concept of “leveraging prior data to enhance the efficiency of BO” as meta-BO. One common meta-BO approach is constructing a GP model to accumulate experiences across all the previous tasks (here: end-users), typically incorporating techniques such as multi-task GP, sparse GP, and other approaches  (Bardenet et al., 2013; Bonilla et al., 2007; Swersky et al., 2013; Yogatama and Mann, 2014). However, the key bottleneck in this class of methods is scalability: GP model fitting scales cubically with the number of observations, i.e., $\mathcal{O}(n^{3})$, where $n$ is the total number of data points. The second approach is called the weighted-sum method; it stores the experience from each previous task as an individual model. When deploying on the new task, these prior models jointly guide the ongoing optimization, where their importance is determined by weights. This method offers slightly better scalability (Li et al., 2022; Schilling et al., 2016; Wistuba et al., 2018), as its computation time only increases linearly with the number of prior models. Some recent works extended this approach for wrist-based input techniques (Liao et al., 2024a) and visual appearance design (Li et al., 2025). However, this approach is still not truly scalable due to the linearly increasing complexity. Notably, a key advantage of HOMI is the unlimited synthetic data generated by the user models. Following the previous approaches, regardless of a single model or a series of models, will not fully unlock the potential of unlimited dataset. To overcome this limitation, our approach follows a third direction – replacing hand-crafted components of BO with deep neural networks, enabling scalable training with large datasets. For instance, a deep kernel (Wistuba and Grabocka, 2021) or fully replacing the surrogate and acquisition function with learned models (Hsieh et al., 2021). Our NAF+ is built upon Volpp et al. (2019), which replaces the manually crafted acquisition function with a deep-neural net, called Neural Acquisition Function (NAF). Pre-trained with data of similar tasks, NAF acquired the specific strategy to pick the next design candidates to be evaluated of this type of tasks. In our context of human-in-the-loop optimization, we define each task as optimizing for a specific user (or synthetic user) under a shared interaction or application. While each user may exhibit different behavior or preferences, we assume there exist shared patterns across users that can be learned and transferred via meta-learning. This formulation is particularly well-suited to our framework, HOMI, where synthetic users (generated from established user models) enable the creation of unlimited training data, a key advantage over traditional approaches. Note that the neural acquisition function in NAF+ is trained via reinforcement learning, which we describe in more detail in the next section.

### v2

# Efficient Human-in-the-Loop Optimization via Priors Learned from User Models

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791976isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interaction techniquesccs: Computing methodologies Machine learning
Figure 1. The general steps of our HOMI framework. Step 1: Model selection — Designers or developers select a model, parameterized by $\theta$, that is relevant to the target task, and optionally fit the model’s parameters (e.g., Fitts’ law, with parameters $a$ and $b$) using minimal user data to better reflect the target context. Step 2: Synthetic user generation — A diverse set of synthetic users is created by sampling different parameter settings $\theta_{1},\theta_{2},\theta_{3},\dots$} from the fitted model’s parameter distribution. Step 3: Meta-BO training — The meta-optimizer interacts extensively with these synthetic users to learn efficient strategies for optimizing across user variability. Step 4: Deployment — The trained meta-BO is deployed with real users and quickly adapts to individual performance or preferences using the learned prior experience.
@@ 1. Introduction
Efficient and effective interface optimization for different users and contexts is a long-standing goal in human-computer interaction (HCI) (Norcio and Stanley, 1989; Zhai et al., 2002; Rao, 2019; Lavie and Meyer, 2010). Model-based optimization approaches address this goal by employing computational optimizers to search for interface designs that maximize predicted user performance, based on established behavioral models. However, most such models capture only population-level behavior, limiting their ability to support personalized interface optimization. Although recent work has attempted to optimize interfaces for individuals or contexts by inferring user-specific or context-specific model parameters during use, this inference is often challenging or time-consuming (Beaumont et al., 2002; Moon et al., 2023), making it difficult to scale or deploy in real-time settings. In recent years, human-in-the-loop optimization (HILO) emerged as a more practical alternative (Lee et al., 2020; Khajah et al., 2016; Chiu et al., 2020; Kim et al., 2017). Unlike model-based approaches that rely on user models, HILO operates directly on actual users: the optimizer proposes a design candidate, users interact with it, their performance or feedback is evaluated, and the optimizer iteratively suggests new designs to improve the user experience. A key advantage of this paradigm is that optimization is driven by users’ real-time evaluations, enabling personalization without requiring a user-model inference step. Various computational techniques, such as Bayesian optimization (Koyama et al., 2017; Liao et al., 2023; Brochu et al., 2010; Dudley et al., 2019), evolutionary and genetic algorithms (Battiti and Passerini, 2010; Meng and Wu, 2008), and heuristic-based techniques (Lee et al., 2020) have been explored to realize HILO. Among these, Bayesian optimization (BO) stands out due to its better performance and minimal assumptions about the task (Frazier, 2018). HILO with BO has been applied in domains such as novel interaction techniques (Liao et al., 2020), design tools (Khajah et al., 2016; Brochu et al., 2010), and wearable interfaces (Liao et al., 2024a; Kim et al., 2017).
While effective for personalizing designs, HILO has a key limitation: the optimization process is often lengthy, typically requiring many rounds of user evaluations to identify a promising interface design (Wang et al., 2022; Pu et al., 2024). The underlying challenge is that the optimizer usually begins with little or no task-specific information, requiring it to explore the design space with random design candidates before gradually converging toward an optimal solution. Such prolonged adaptation not only exposes users to suboptimal designs, but also forces them to repeatedly adjust their behavior, or even re-learn interaction strategies as the interface continues to change. This is especially problematic for interactions that demand instant and stable performance, such as most input techniques. Ideally, an optimizer should therefore identify a promising, or even optimal, design within as few trials as possible, and then fix the interface to provide a stable experience for the remaining use. Consequently, efficiently finding high-performing designs early has become a central goal in HILO and broad adaptive interfaces (Liao et al., 2024a; Koyama et al., 2017).
@@
Recent work has attempted to enhance the adaptation efficiency by incorporating prior data from past users to initialize or guide the optimization process, a concept known as transfer learning or meta-learning. For example, previous works have explored constructing a single surrogate model based on all the previous users’ evaluation results (Liao et al., 2024b), storing a series of separate prior user models and aggregated them via a weighted-sum approach during deployment (Liao et al., 2024a), or maintaining an adaptive surrogate model using Bayesian neural networks to accumulate growing data (Liao et al., 2025). These works demonstrated that augmenting HILO with prior experience is a viable and general path toward solving the cold-start issue. However, a key bottleneck remains: the reliance on real human data. Prior transfer- or meta-learning–based methods require data from at least ten users to be effective (Liao et al., 2024b, a), which can be prohibitively expensive or impractical. Moreover, any changes to the interaction task or interface require fresh rounds of data collection, severely limiting scalability. This motivates a fundamental question: Is it possible to train computational optimizers for higher online efficiency, yet, without relying on real human data?
@@
In this work, we propose a new computational framework called Human-in-the-Loop Optimization with Model-Informed Priors (HOMI), illustrated in Figure 1. Here, “model-informed priors” refer to the learned optimization strategies acquired by training with predictive models. Rather than relying on real user data for training the optimizer, HOMI leverages data from synthetic users simulated from these models. This offline training phase allows Bayesian optimization to develop a broad understanding of the design space and to learn effective adaptation strategies across diverse user contexts before online optimization begins. A key advantage of using model-informed priors is that the optimizer can be trained with a wide variety of synthetic users, enabling generalization across different usage scenarios. In contrast, training with real users limits the optimizer’s knowledge to a small number of specific individuals, reducing its adaptability. At deployment, the trained optimizer switches to a human-in-the-loop mode, integrating prior knowledge with live user feedback to efficiently optimize the interface design in real time.
@@
We further demonstrate the effectiveness of our HOMI by applying NAF+ to real-time adaptation of a mid-air keyboard for direct touch input, which is an essential VR text entry technique, in a study with real users. The training involved synthetic users constructed via established models in HCI, including Fitts’ law (MacKenzie, 1992b) and typing error model (Bi and Zhai, 2016). The results showed HOMI enabled by NAF+ outperforms the established baselines, including both manual keyboard setting and standard BO in that it significantly reduced the number of iterations required to personalize the keyboard.
@@ 2. Background & Related work
We review the key topics that are relevant to this work, especially focusing on model-based optimization in HCI, HILO, and the techniques that augment Bayesian optimization with prior data for efficient interface adaptation.
@@ 2.1. Model-based Optimization in HCI
Optimization is an essential process in HCI and design, even if they are not explicitly articulated as such. In traditional user-centered design workflows, optimization is performed manually: designers generate a set of candidate designs, which are then evaluated through user testing and iteration (Abras et al., 2004). The field of computational interaction (Oulasvirta et al., 2018) aimed to integrate computational optimization methods in the design process to automatically identify optimal design candidates. This idea is not new: engineering fields have long used optimization workflows in which a solver proposes candidate designs, and these are evaluated through domain-specific models (e.g., physical, mechanical, chemical) to identify optimal solutions (Martins and Ning, 2021). This approach has proven successful in fields such as mechanical engineering (Rao et al., 2011), chemical engineering (Bhaskar et al., 2000), and architectural design (Michalek et al., 2002). HCI researchers have adopted similar methods: deploying computational solvers over user models to identify optimal interaction designs. Examples include menu optimization (Bailly et al., 2013; Todi et al., 2021), keyboard layout design (Feit et al., 2021; Karrenbauer and Oulasvirta, 2014), and user interface layout adaptation in AR/VR environments (Lindlbauer et al., 2019; Evangelista Belo et al., 2022), and more.
With the rapid growth of user modeling approaches, model-based optimization has shown success in generating promising design candidates. A key opportunity can extend to enabling adaptive interfaces: given a new user and context, in principle, one could first infer that user’s model parameters (in other words, inverse modeling or model inference), then run a computational solver over that personalized model to derive an optimized interface. However, this ideal pipeline is rarely feasible in practice. Inverse modeling is computationally expensive and often ill-posed (Kangasrääsiö et al., 2017; Moon et al., 2022): multiple parameter combinations can reproduce the same behavioral observations. As a result, most model-based optimization in HCI has targeted an abstract average user, or a fixed user group (Sarcar et al., 2018; Todi et al., 2021), and the deployed systems typically do not update or adapt their underlying model based on live observations.
This paper circumvents this inverse-modeling bottleneck. Instead of inferring exact model parameters for each user, we pursue human-in-the-loop adaptation to directly identify the optimal interfaces for each individual with minimal user intervention. The HOMI framework builds on the foundation of model-based optimization, but reframes user models not as fixed targets to optimize against, but as generators of synthetic user data for training an optimizer.
@@ 2.2. Human-in-the-Loop Bayesian Optimization
Human-in-the-loop optimization (HILO) is an emerging optimization paradigm in HCI that does not rely on predefined user models. Instead, HILO learns and optimizes directly from online observations, which may include user performance, preferences, or other forms of interaction feedback. The objective function $f$ is usually a user-centered metric (e.g., completion time (Liao et al., 2025), preferences (Koyama et al., 2020), etc.) that we want to optimize for (Liao et al., 2020, 2024a; Kim et al., 2017; Yamamoto et al., 2022). Each unique design candidate $x$ within the designated design space $\mathcal{X}$ results in a user-related objective function value, $f(x)$. The goal of HILO is to identify the global optimal design $x^{*}$ that maximizes (or minimizes) the objective value $f(x^{*})$. Such optimization problems are challenging since the only way to access the user-relevant objective function is through costly user evaluations. Bayesian optimization (BO) is a popular computational method for human-in-the-loop optimization due to its higher sample-efficiency than other approaches (requiring fewer user evaluations), minimum requirements and assumptions of the problem, and reliable performances (Mockus, 1975; Shahriari et al., 2016). Within Bayesian optimization, a surrogate model, typically a Gaussian Process regression (GP), captures the properties of the target function. This model is continually refined with online function observations (the result of user evaluation), enhancing its prediction accuracy. An acquisition function utilizes the information of the surrogate model to compute the acquisition value (i.e., the worth or potential value) of a new candidate design. The candidate with the highest acquisition value is then selected for user evaluation in the subsequent iteration.
@@
Bayesian optimization has found many human-involved applications where the sample efficiency is particularly important, such as exoskeleton systems (Kim et al., 2017, 2019; Tucker et al., 2020), wearable interfaces (Liao et al., 2023; Guo et al., 2019), haptic displays (Liao et al., 2023; Catkin and Patoglu, 2023), input techniques (Liao et al., 2024b), and design tools (Brochu et al., 2010; Koyama et al., 2017; Mo et al., 2024; Langerak et al., 2026). Despite its robust performance and relatively better efficiency over other optimization approaches (simulated annealing, evolutionary algorithms, etc.), a key limitation of BO lies in its “cold-start” problem. Since BO generally starts without any prior knowledge of the objective function, it must rely on random exploration in the early stages, which can slow down convergence. To address this, recent work has explored incorporating transfer learning and meta-learning to accelerate the optimization process, which we will review in the next subsection.
2.3. Meta-learning for BO
Meta-learning (Wang, 2021) is a machine learning paradigm that aims to accelerate learning on a novel task by leveraging experience from previous tasks with a similar structure. This idea has also been extended to Bayesian Optimization (BO), where it is known as meta-BO: instead of optimizing each problem from scratch, the optimizer draws on past optimization tasks to guide search on a new one. In our setting, a task corresponds to optimizing the interface for a specific user, making meta-BO a natural fit, as it enables rapid adaptation when a new user arrives. Recent surveys, such as Bai et al. (2023) provides a comprehensive overview of recent meta-BO methods. A common approach is to merge all data from prior tasks into a single unified surrogate model, often implemented with multi-task Gaussian Processes (GPs), sparse GPs, or other implementations (Bardenet et al., 2013; Bonilla et al., 2007; Swersky et al., 2013; Yogatama and Mann, 2014). This unified surrogate is then used to guide optimization for a new user. Another line of work constructs separate surrogate models for each task, and combines their predictions through a weighted sum for a new task (Li et al., 2022; Schilling et al., 2016; Wistuba et al., 2018). These methods are conceptually simple, easy to implement, and have been applied in HCI contexts (Liao et al., 2024a; Li et al., 2025). However, both suffer from scalability issues as the number of past tasks grows: the unified model requires cubic-time GP inference, whereas the weighted-sum methods scale linearly.
In this work, we exploit the fact that user models can generate unlimited synthetic data, providing far richer information than prior data in typical meta-BO settings. However, leveraging this large-scale dataset requires meta-BO methods that scale with the number of tasks and volume of data, which classical approaches cannot satisfy. We therefore follow a third direction in meta-BO: replacing hand-crafted components of BO with deep neural networks (Wistuba and Grabocka, 2021; Hsieh et al., 2021), enabling efficient inference even with a large number of prior tasks. Our NAF+ is built upon Volpp et al. (2019), which replaces the manually crafted acquisition function with a deep neural network, called Neural Acquisition Function (NAF). Pre-trained with data from optimization across users, NAF learns a general strategy to select the next design candidates promising to be evaluated. While individual users may differ in their behaviors or preferences to some extent, we assume an underlying shared structure across user-optimization tasks that meta-BO can learn and transfer. Note that the acquisition function in NAF+ is trained via reinforcement learning, which we describe in more detail in the next section.